### v2

# Surface bubble lifetime in the presence of a turbulent air flow, and the effect of surface layer renewal

Abstract
Surface bubbles in the ocean are critical in moderating several fluxes between the atmosphere and the ocean. In this paper, we experimentally investigate the drainage and lifetime of surface bubbles in solutions containing surfactants and salts, subjected to turbulence in the air surrounding them modelling the wind in the ocean. We carefully construct a setup allowing us to repeatably measure the mean lifetime of a series of surface bubbles, while varying the solution and the wind speed or humidity of the air. To that end, we show that renewing the surface layer is critical to avoid a change of the physical properties of the interface. We show that the drainage of the bubbles is well modelled by taking into account the outwards viscous flow and convective evaporation. The lifetime of surface bubbles in solutions containing no salt is given by the time when evaporation becomes the main draining mechanism and independent on surfactant concentration. When salt is added, the same scaling is valid only at high surfactant concentrations. At low concentrations, the lifetime is always smaller and independent of wind speed, owing to the presence of impurities triggering a thick bursting event.
1 Introduction
Bubbles at the surface of the ocean play a key role in atmosphere-ocean interactions. When they burst, they release droplets into the air that facilitate heat, momentum and mass exchanges (Deike,, 2022). These droplets carry with them the chemical and biological contents of the top layer of the ocean: salts but also any organic or artificial contaminants that may be present (Cunliffe et al.,, 2013; Burrows et al.,, 2014; Veron,, 2015). The smallest droplets may be entrained in upper layers of the atmosphere where they serve as an important source of cloud condensation nuclei (Lewis and Schwartz,, 2004). Understanding and being able to predict the number and size of these droplets is therefore important to better model atmospheric processes.
There are several pathways for the generation of droplets from bubble bursting (Lewis and Schwartz,, 2004; Jiang et al.,, 2022; Deike et al.,, 2022; Jiang et al.,, 2024), but the one we are most interested in is the ”film drops” mechanism as it may be able to produce very small droplets associated with a large uncertainty (Dubitsky et al.,, 2023; Zinke et al.,, 2022). Film drops are generated during the retraction of the bubble cap, through a centrifugal instability (Blanchard and Syzdek,, 1988; Lhuissier and Villermaux,, 2012; Jiang et al.,, 2022) and are typically of similar size as the thickness of the cap from which they were generated. It is therefore important to know the drainage rate and lifetime of those surface bubbles in order to know the size of droplets generated. The drainage of surface bubbles has been thoroughly investigated by many lab studies in various conditions (Lhuissier and Villermaux,, 2012; Champougny et al.,, 2016; Poulain and Bourouiba,, 2018; Poulain et al.,, 2018; Pasquet et al.,, 2022; Shaw and Deike,, 2024). In the conditions relevant to oceanic surface bubbles, the drainage is driven by a curvature induced pressure gradient, resulting in the thickness of the bubble cap $h$ decreasing like $t^{-2/3}$, until at long times (a few tens of seconds) direct evaporation from the cap dominates, and the thickness decreases linearly in time. Corrections can be added to this base flow to account for surface tension gradients that can form because of evaporation from the top of the bubble (Chandran Suja et al.,, 2018; Poulain et al.,, 2018; Shaw and Deike,, 2024).
The lifetime of surface bubbles ranges between a few to tens of seconds, but precisely predicting the expected lifetime for a given set of physical parameters is challenging. Previous authors have noted the difficulty in even measuring lifetime as it is extremely sensitive to the type and amount of surfactants present in the solution and can vary drastically without any change of the physical parameters of the study (Poulain et al.,, 2018). The maximal lifetime that a bubble can live is set by evaporation as the fast removal of fluid from the cap at late times make it highly susceptible to bursting (Miguet et al.,, 2020; Roux et al.,, 2022). The difficulty in accurately predicting lifetime lies in the identification of a single bursting criterion: previous studies have shown that bubbles can burst at a wide range of thicknesses, some much larger than the thickness at which Van der Waals forces act. One possible explanation for bursting events at large thickness is proposed in Néel and Villermaux, (2018) or Poulain et al., (2018) as the formation of a hole in the cap due to sufficiently large local surface tension gradients caused by surfactants or impurities at the interface.
This study aims towards measuring mean lifetimes and drainage curves of surface bubbles in realistic oceanic conditions. To this end, we use solutions representative of ocean water and introduce agitation in the air above the water surface. Indeed, the studies mentioned above investigated the behavior of bubbles in laboratory settings, with the water surface flat and motionless, but bubbles at the surface of the ocean are constantly pushed by the motion of the ocean surface as well as subjected to turbulence in the air from the atmospheric boundary layer. This turbulence potentially having the effect of altering drainage through bubble cap entrainment or altering the evaporative rate of bubbles. In addition, we recently showed that mixing in the air around thin films have the potential to drastically affect lifetime and reduce the large fluctuations observed in previous studies (Aurégan and Deike,, 2025).
The capacity of oceanic aerosols to serve as cloud condensation nuclei strongly depends on their chemical composition (Burrows et al.,, 2014). As a result, the exact composition of the solution in ions, surfactants, but also organic material may affect the final coupling between atmosphere and ocean. In this study, we use model solutions that can readily be prepared in the lab: containing either only NaCl in seawater concentration or a mixture of salts representative of those that can be found in the ocean. We use Sodium Dodecyl Sulfate (SDS) as a surfactant as it has often been used in the atmospheric aerosols’ literature to model the naturally occurring lipids in the ocean (Li et al.,, 1998; Tuckermann,, 2007; Burrows et al.,, 2014). Other compounds such as palmitic or stearic acid are though to better model the oceanic surface layer (Burrows et al.,, 2014), but SDS has the advantage of being easily available and a wide body of literature on its physical properties (Fainerman et al.,, 2010). Finally, our solutions contain no organic macromolecules (polysaccharides or proteins) that are sometimes used in atmospheric chemistry studies.
This paper is organized as follows: in Sec. 2 we present the experimental setup and in Sec 3 the steps that have been taken to ensure the repeatability of the lifetime measurement with surface layer renewal. Then in Sec. 4 we present mean lifetime and drainage data for solutions containing only surfactants and surfactants and salts. Finally, in the discussion (Sec. 5) we analyze the differences between those cases and propose a rationalization. In the following, we make the distinction between different classes of surface active compounds. To avoid confusions, we define here different terms for these classes (similarly to Mcnair et al., (2025)): surfactants refer to surface active molecules that are commonly used detergents and that we have purposefully added to the solution. Co-surfactants are similar molecules that can be found in trace amounts in commercially available surfactants powders. Finally, impurities are compounds that have an influence on surface properties that can come from various sources in the environment (dust, tap water organics, etc.) and that we have not purposefully added. Adding salts has both the effect of introducing ions in the solution that may influence the surface properties through electrostatic interactions and also adds impurities that are present in the powder that we use.
2 Material and Methods
Figure 1: Schematic of the experimental setup. Bubbles are injected underwater through a needle connected to a syringe pump. Bubbles rise to the surface and are monitored via a top-down viewing camera and backlight system. Turbulence in the air is generated via two counter-rotating fans and the humidity of the air is also controlled. Finally, a peristaltic pump continuously overflows the container where bubbles are generated into a larger container, avoiding surface ageing.
We measure the lifetime of surface bubbles in various conditions. The bubbles are injected 5 cm below the surface and rise freely up to the surface (70 mm $\times$ 70 mm) where they can move with no restrictions. The water level is slightly above the edge of the container such that the bubbles are naturally pushed away from the edges. The container is continuously overflowed into a larger reservoir using a peristaltic pump such that the surface layer is renewed constantly (detailed below). We record the bubbles with a camera (Basler acA1920-40um) placed above the surface, allowing us to measure the radius of the bubbles as well as their lifetime (from the moment they reach the surface to the final bursting event). A live detection algorithm detects the presence of a bubble on the surface from the camera feed and allows us to send a new bubble only when the previous one has burst. We run this experiment for a large number of bubbles, up until the relative uncertainty on the mean lifetime is below 5% or at least 150 repetitions are reached. Note that while this is enough to accurately measure averaged quantities, it does not allow us to precisely assess the lifetime distribution as we focus here on obtaining data for a large range of conditions.
The solutions used to generate bubbles are mixes of surfactant (Sodium Dodecyl Sulfate or SDS, with a concentration $c_{\textsc{SDS}}$) and salt (either pure NaCl or artificial sea salt (ASTM D 1141-98)). The conditions explored in this study are summed up in Tab. 1. In addition to sodium and chloride ions, the sea salt solutions contain magnesium, calcium, sulfate and potassium ions in significant proportions. A small amount of the initial sea salt powder does not dissolve and remains as floating particles in the solution. SDS is a negatively charged surfactant and as a result, the combined effect on surface tension or other surface properties of a small amount of SDS and a small amount of ions in solution can be very different from each of the components alone. We measured the Gibbs isotherms of all the solutions used in this study in a Langmuir trough using the Wilhemy plate method. These isotherms can be found in Appendix A.
The bubbles are generated by pushing air through a needle tip (inner diameter 1.6 mm) using a syringe pump, allowing us to obtain bubbles of radius $R=1.6\pm 0.1$ mm. Because of the large variations in surface tension, there is a slight trend towards smaller radii at low surface tension (except for the highest SDS concentration with sea salt where we decided to increase the syringe volume to obtain a radius closer to the mean value). The effect of the slight radius changes are taken into account through the drainage timescale ($\tau_{d}$, defined in Eq. (4)).
Solution No salt 35 g/L NaCl Sea salt
$c_{\textsc{SDS}}$ (µmol/L) $10$ $200$ $3000$ 101010104 $10$ $50$ $100$ $200$ $400$ $10$ $20$ $50$ $100$ $200$ $400$
$\gamma_{0}$ (mN/m) $69.9$ $67.0$ $47.6$ $35.6$ $62.8$ $47.0$ $39.1$ $35.2$ $20.0$ $54.4$ $50.5$ $43.3$ $38.2$ $31.7$ $28.3$
$R$ (mm) $1.7$ $1.7$ $1.5$ $1.4$ $1.6$ $1.5$ $1.4$ $1.4$ $1.4$ $1.6$ $1.5$ $1.5$ $1.5$ $1.3$ $1.7$
$\tau_{d}$ (ms) $47.9$ $52.7$ $48.3$ $56.1$ $42.9$ $49.7$ $50.6$ $55.5$ $83.2$ $53.0$ $50.0$ $55.5$ $55.3$ $51.9$ $137.7$
Table 1: Physical parameters of the solutions used in this study: SDS concentration $c_{\textsc{SDS}}$, equilibrium surface tension $\gamma_{0}$ (measured using the Langmuir trough at zero compression), average radius of the bubbles generated $R$, and associated drainage timescale $\tau_{d}$ (defined in Eq. (4))
The setup is placed in a container where the air is monitored. The humidity is duty-cycled controlled using a PID controller alternatively injecting dry or humid air in the chamber. This setup has been previously detailed in Boulogne, (2019). We create a turbulent flow above the surface using two counter-rotating fans (5 cm radius) placed 17 cm apart on each side of the container (see Fig 1). The fans have four 3D printed blades and their angular velocity can be modulated. Throughout the study we use five conditions: 0, 50, 275, 460, and 610 rotations per minute.
[…]
Figure 2: (a) Average flow velocity $\mathbf{U}\right>$ above the bath surface. (b) Root mean square velocity $\mathbf{u}^{\prime}$ above the bath surface. The withe area at the bottom is the location of the container, the black circle shows the typical size of bubble in this study ($R=1.5$ mm). Both of these snapshots correspond to the case where the angular velocity of the fan is 460 RPM.
The integral length scale $L_{\rm int}$ is computed by integrating the longitudinal autocorrelation of the velocity fluctuations and the dissipation rate $\varepsilon$ is estimated via $\varepsilon=0.7\left|\mathbf{u}^{\prime}\right|^{3}/L_{\rm int}$ (Vassilicos,, 2015). Finally, the integral Reynolds number is $Re=\left|\mathbf{u}^{\prime}\right|L_{\rm int}/\nu_{a}$, with $\nu_{a}$ the kinematic viscosity of air. The turbulent properties are summed up in Tab. 2, the values reported are the averages over the 5 cm above the surface of the bath
[…]
0 0.0 0.0 2.3 0.0 0
50 2.0 2.0 1.2 $4.0\times 10^{-4}$ 16
275 9.0 13.5 1.6 0.14 142
460 15.3 24.1 2.0 0.50 315
610 22.4 33.7 2.1 1.59 466
Table 2: Average values of the turbulence parameters.
3 Surface ageing and the need for overflow
Figure 3: Effects of surface ageing on bubble lifetime. (a) and (c) show the distribution of lifetimes for several consecutive runs of 150 bubbles. The shaded area illustrates the whole distribution while the plus signs show the mean value. In (a) there is no overflow but after 4 hours (vertical red line) the surface layer is removed with a syringe ($c_{\rm SDS}=200$ µmol/L, $\mathcal{R}_{H}=50\%$, no wind). In (c) the overflow is continuously turned on ($c_{\rm SDS}=200$ µmol/L, $\mathcal{R}_{H}=50\%$, no wind, sea salt). (b) Schematic of the adsorption curve for SDS displaying two typical timescales: at early times SDS is adsorbing to the interface (orange area) and at late times the surface ages and the surface tension slowly decreases (red). (d) Measurements of the surface tension at long times with (green) and without (blue) continuous overflow of the solution. The surface layer is ageing and surface tension decreases in the case without overflow while it remains constant when the overflow is activated. (e) Lifetime as a function of the Reynolds number in the high concentration ($c_{SDS}\gtrsim$ CMC) limit where the effects of surface ageing are the least visible. The blue curve is the data with no overflow, while the green curve is the data with overflow. Vertical bars represent one standard deviation.
Surface properties need particular attention to be able to study bubble lifetime. Indeed, our goal in this study is to investigate the effect of various parameters on the mean lifetime of surface bubbles. In order to have robust conclusions however, we need to make sure that we can repeatably measure the mean lifetime for a given set of parameters. In particular, repeating the same experiment should yield the same result within some margin of error. This is not trivially achieved with surface bubbles as was previously discussed by Poulain et al., (2018). It is not obvious in our case as well and illustrated in Fig. 3 (a) (using a solution containing 200 µmol/L of SDS and no salt): while measuring the lifetime of 150 bubbles gives a good measurement of the mean (a shaded area illustrates the distribution for a group of 150 bubbles and a plus sign indicates the mean), repeating the same 150 bubble run reveals a slow drift of the mean lifetime towards short times. This drift induces lifetime differences that can be of tens of seconds, which is of the same order of magnitude as the differences we expect to see from changing the wind speed or the solution. Humidity and fan speed are continuously recorded and do not vary in time, temperature is also regularly checked and does not significantly vary.
Our interpretation of this drift is the ageing of the surface layer and therefore a modification of its physical properties. This is made evident by the fact that, after four hours (or five runs) we removed the surface layer of the solution with a syringe (vertical red line). This removal results in a very clear change of the distribution, which becomes similar to the one of the first run. The drift towards short lifetime then resumes. The ageing of the surface can be caused by several effects: first, surface active molecules can be present in the air surrounding the setup in the experiment and slowly fall onto the surface (even though in our case the setup is in a closed box and the air filtered before entering). Similarly, surface active molecules can be found in small amounts because of imperfect cleaning of the experimental setup. Because of the steps taken to prevent these sources of impurities, we believe that these effects are small. Finally, it has been show previously that even high purity surfactant products often contain remnants of other molecules used in the fabrication process that can have surface active properties such as dodecanol in the case of SDS (Fainerman et al.,, 2010; Miguet et al.,, 2020). These co-surfactants very slowly adsorb to the interface (over several hours) and strongly reduce the surface tension of the solution. One possibility is then to purify the surfactant used (as performed by Miguet et al., (2020)) to remove the traces of other surfactants, however SDS also slowly undergoes hydrolysis to naturally form dodecanol, and it is likely that completely getting rid of co-surfactant from other sources would be difficult.
Instead, we take advantage of large difference in timescale between the adsorption of SDS and all other co-surfactants. A sketch of the typical surface tension over time of an SDS solution (made without purifying the surfactant) can be found in Fig. 3 (b) based on measurements of the same curve in Fainerman et al., (2010). Depending on the concentration, the SDS adsorbs to the surface in a typical timescale of milliseconds to at most tenths of seconds (orange area), after that the surface tension is relatively constant until the co-surfactants have significantly adsorbed to the interface, typically in thousands of seconds (red area). By continuously renewing the surface we can therefore obtain an interface that remains in this intermediate range where surface tension can be considered constant. Any new surface element quickly equilibrates with the rest of the surface and is then slowly advected towards a designated location on the edge of the tank, therefore not ageing for more than a minute. Experimentally, this is done by continuously overflowing the container where bubbles are generated into a larger container below using a small peristaltic pump (see Fig. 1).
We have measured the static surface tension of SDS solutions using the Wilhelmy plate method with and without overflowing for timescales larger than 10 seconds and see that the drift is indeed prevented by overflowing the solution (shown in Fig 3 (d)). The result of implementing this approach in our setup can be seen in Fig 3 (c) where we repeat runs of 150 bubbles as in (a) but with the overflow activated. In contrast to the previous case, the lifetime distribution remains statistically similar for several hours. It should be noted that we have checked that the pump flow rate (and therefore the renewal rate of surface elements) does not affect the lifetime measurably and that this result is not specific to the SDS concentration (200 µmol/L) used here. Regular renewal of the surface layer, in opposition to a still interface over a motionless subphase is also a reasonable working assumption when considering bubbles in the ocean or in rivers where there is turbulence close to the surface.
At high concentrations (around the Critical Micelle Concentration or CMC, and static surface tensions of about 30 mN/m), the effect of surface ageing is less visible because the effects of SDS are already very important. This allows us to compare lifetimes obtained in the cases with or without overflow (Fig 3 (e)). In this limit, both approaches yield similar results, validating the use of overflow to study the lifetime of surface bubbles. In practice, we never use the same solution for more than two consecutive days, and we estimate the minimal uncertainty on the mean lifetime measurement to be about one second.
4 Results
Figure 4: (a) Illustrations of thickness variations in the bubble cap using color interferometry. From left to right the Reynolds number is 0, 142, and 466. (b) Lifetime as a function of the integral Reynolds number. Each circle represents a single bubble lifetime, the boxplot adds additional statistical information: (green line) median, (orange dashed line) mean, (box width) interquartile range, (whiskers) furthest data point within 1.5 times the interquartile range from the box. The data corresponding to a fan RPM of 50 or $Re=18$ has been slightly shifted for clarity. The data presented is obtained with 200 $\mu$mol/L of SDS and at a relative humidity of 50 %. (b) Mean lifetime as a function of the parameter for several surfactant concentrations, wind speeds and relative humidities. Color indicates SDS concentration and markers show relative humidity. The dashed line is a power law of slope 0.5.
4.1 Qualitative observations
The presence of wind strongly disrupts the classical picture of the motion of the patches in the bubble cap (Illustrated with three panels of Fig. 4 (a)). In these frames white light creates interference pattern within the bubble cap and provides a view of thickness variations. In the top panel (no wind) we see patches generated at the foot of the bubble through marginal regeneration and rising in the cap. These patches allow the exchange of fluid between the bubble cap and the bath and are responsible for the drainage of the cap. They are generated through an instability of the pinch at the foot of the bubble (Lhuissier and Villermaux,, 2012; Gros et al.,, 2021; Trégouët and Cantat,, 2021). These patches are thinner than the rest of the film and therefore buoyant, and they detach from the pinch at the foot when they are large enough. Each patch generated therefore corresponds to a loss of volume in the bubble cap, and it has been shown that the volume change due to generation of these patches exactly matches the viscous limited flow out of the cap (Miguet et al.,, 2021). The motion of these patches in the cap can also be taken into account to model flows induced by surface tension gradients either due to evaporation or heating for instance (Poulain et al.,, 2018).
This picture is largely modified by the presence of wind (middle and bottom panel). Instead of rising straight up because of buoyancy, the patches are entrained by the wind in the air close to cap. They are therefore constantly being deformed and pushed around by the strongly intermittent wind close to the bubble. In the middle panel we can still discern the patches, but some of them have a mostly horizontal velocity instead of vertical and their shape is quite different from the classical picture. In the bottom panel (largest wind forcing) the cap is almost completely mixed with the patches now taking the form of horizontal bands in the cap. In the case with the lowest wind forcing ($Re=16$, not shown), the motion of the patches varies between being very similar to the no wind case and being entrained by the wind as an eddy reaches the proximity of the bath surface. We have verified that the presence or absence of the overflow has no significant effect on the bubble cap images or the conclusions above.
These observations could have important consequences for the drainage of the bubble as the size selection of the patches occurring at the foot of the bubble is most likely dominated by wind forcing instead of buoyancy for stronger wind speeds. As the patches are setting the drainage rate of the bubble, this could strongly alter the drainage profiles. However, as we show later in Fig. 5, the effect of the wind is important on the evaporative rate but not on marginal regeneration or capillary pressure induced drainage.
Wind strongly reduces the mean lifetime of surface bubbles in solutions of pure surfactants. For a given solution and relative humidity, the lifetime can vary by a factor up to 5 only due to the intensity of the wind forcing. Fig. 4 (b) illustrates this effect, (with 200 µmol/L of SDS and a relative humidity of 50%) where we can see several effects of the presence of wind: each point is an individual bubble lifetime measurement and the box plot shows the most important metrics of the distribution. The mean lifetime (orange horizontal lines) monotonously decreases as the Reynolds number increases, and so does the spread around the mean. The reduction of the spread is reminiscent of our previous study on soap films (Aurégan and Deike,, 2025), where we found that mixing the air avoids unwanted local fluctuations of the humidity field close to the film. A similar process may be happening here as conditions get more homogeneous and statistically invariant in time as the wind forcing is increased. The reduction in the mean lifetime may a priori be due to many effects: early bursting due to a perturbation from the wind, increased drainage through direct evaporation from the cap or increased flow through the bubble foot because of the strong cap mixing for instance. In the following we derive a drainage model taking into account only direct evaporation and show that it is sufficient to explain most of the observed variation of lifetime.
4.2 Drainage
We model the effect of the wind on bubble lifetime by taking into account its effect on the evaporative rate. The effect of the wind-induced motion in the cap is not considered in the following. There are two mechanisms responsible for the drainage of the cap of the bubbles: viscosity limited flow from the cap towards the bath due to the capillary pressure through the foot of the bubble and direct evaporation from the cap. Following the derivation from previous studies (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018), the velocity through the foot of the bubble reads $u=\alpha_{d}(\gamma/\mu)(h/R)^{3/2}$, with $\gamma$ and $\mu$ the surface tension and viscosity of the liquid, respectively. $\alpha_{d}$ is non-dimensional a constant of order unity. The evaporative flux for an arbitrary convection in the air can be written as
$j_{e}=\alpha_{e}Sh\frac{Dc_{\rm sat}(1-\mathcal{R}_{H})}{R\rho},$ (1)
where $D$ is the diffusivity of water vapor in air, $c_{\rm sat}$ the mass saturation concentration, $\rho$ is the density of the liquid, and $\alpha_{e}$ a non-dimensional constant. Finally, $Sh$ is the Sherwood number, quantifying the importance of convective evaporation compared to diffusion alone, which will vary depending on the regime of interest. Mass conservation in the cap of the bubble reads:
$S\frac{\mathrm{d}h}{\mathrm{d}t}+Phu+Sj_{e}=0,$ (2)
with $S$ and $P$, the surface and perimeter of the bubble, respectively. Substituting the values of $u$ and $j_{e}$ and using $P/S\sim\ell_{c}/R^{2}$, with $\ell_{c}=\sqrt{\gamma/\rho g}$ the capillary length (Lhuissier and Villermaux,, 2012), we obtain the equation describing the drainage of the cap:
$\frac{\mathrm{d}h}{\mathrm{d}t}+\alpha_{d}\frac{\gamma\ell_{c}}{\mu}\frac{h^{5/2}}{R^{7/2}}+\alpha_{e}Sh\frac{Dc_{\rm sat}\left(1-\mathcal{R}_{H}\right)}{R\rho}=0.$ (3)
We normalize the thickness in the previous equation using the estimated thickness at generation $h_{0}$, and obtain two characteristic timescales $\tau_{d}$ and $\tau_{e}$ representing drainage and evaporation, respectively:
$\tau_{d}=\frac{R^{7/2}\mu}{\gamma\ell_{c}h_{0}^{3/2}}\,{\rm and}\,\tau_{e}=\frac{R\rho h_{0}}{Dc_{\rm sat}},$ (4)
and the drainage equation in dimensionless variables ($\tilde{h}=h/h_{0}$ and $\tilde{t}=t/\tau_{d}$) reads:
$\frac{\mathrm{d}\tilde{h}}{\mathrm{d}\tilde{t}}+\alpha_{d}\tilde{h}^{5/2}+\alpha_{e}\frac{\tau_{d}}{\tau_{e}}Sh\left(1-\mathcal{R}_{H}\right)=0.$ (5)
This model is similar to the one employed by Poulain and Bourouiba, (2018) with the difference that we explicitly specify the evaporative rate in several conditions. The value of the Sherwood number depends on the Reynolds number in the air and the relative humidity. It can be divided into natural and forced convection depending on the value of the Reynolds number. At low Reynolds numbers (i.e. the two lower wind cases of our study), the evaporation is dominated by natural convection. In that case, the evaporation of the surface of the bath creates a plume of lighter high humidity air that rises in the center of the bath. The relevant dimensionless parameter in this problem is the Grashof number comparing evaporation-induced buoyancy and viscous dissipation: $Gr=\Delta\rho_{a}/\rho_{a}\times g\mathcal{L}^{3}/\nu_{a}^{2}$, with $\rho_{a}(\mathcal{R}_{H})$ the density of the air, $\mathcal{L}$ a typical length scale of the bath surface and $\nu_{a}$ the kinematic viscosity of air. The Grashof number can be simplified into $Gr\approx(1-\mathcal{R}_{H})Gr_{0}$ with $Gr_{0}$ the maximal value of the Grashof number, occurring with a perfectly dry air in the chamber (Aurégan and Deike,, 2025). In our setup the Grashof number is typically $10^{4}$. (Dollet and Boulogne,, 2017) have measured a scaling for the Sherwood number associated with the evaporation of the flat surface of the bath in this situation:
$Sh_{N}=\frac{\pi}{2}\left(a_{1}Gr^{0.18}+a_{2}\right),$ (6)
with $a_{1}=0.31$ and $a_{2}=0.48$ two dimensionless constants. Note that the theory in the same reference predicts a scaling with the Grashof number with an exponent $1/5$. In order to use the experimentally fitted values for $a_{1}$ and $a_{2}$, we also use the fitted exponent of $0.18$ (fitted in the case of a circular bath while ours is square). It should be noted that the theory gives a thickness of the boundary layer that is developing above the surface scaling as $\delta_{\mathcal{R}_{H}}\sim R/Gr^{1/5}\approx 5$ mm. The bubbles are therefore fully inside of the concentration boundary layer and expected to evaporate with a similar Sherwood number as the surface of the bath.
At higher Reynolds numbers, the airflow above the bath is turbulent and dominates the evaporation. The rate at which bubbles evaporate is limited by the rate of the renewal of dry air elements close to the surface, which depends on the Reynolds number. The exact form of the relationship between the Sherwood and the Reynolds number depends on the properties of the flow, the geometry of the problem and can also depend on the interfacial properties (Theofanous et al.,, 1976; Turney and Banerjee,, 2013; Wissink et al.,, 2017). This relation can be found by using Higbie penetration theory (Higbie,, 1935; Farsoiya et al.,, 2023): the mass transfer velocity $k_{e}$ is given by $k_{e}\propto\sqrt{D/\tau}$, with $\tau$ the timescale at which air surface elements are renewed. If the relevant timescale for surface renewal is the integral timescale $\tau_{\rm int}=L_{\rm int}/\left|\mathbf{u}^{\prime}\right|$, then the Sherwood number $Sh_{F}=k_{e}L_{\rm int}/D$ reads:
$Sh_{F}\propto Sc^{1/2}Re^{1/2}.$ (7)
If however, the relevant timescale is given by the smallest scales of the turbulent motion, then $\tau\sim(\nu/\varepsilon)^{1/4}$, and we get:
$Sh_{F}\propto Sc^{1/2}Re^{3/4}.$ (8)
In the following, we use a Sherwood number combining the effects of natural and forced evaporation. The forced convection scaling based on the integral timescale is expected to be valid at low Reynolds numbers, while the one based on the small turbulent scales is expected to be valid at higher Reynolds numbers (Theofanous et al.,, 1976). In the following, we will therefore assume that the relevant timescale is the integral timescale. Finally, the Sherwood number can be written as the sum of a term dependent on $1-\mathcal{R}_{H}$, a constant term and term dependent on the Reynolds number:
$Sh=Sh_{N}+Sh_{F}=\frac{\pi}{2}\left(a_{1}Gr^{0.18}+a_{2}\right)+Sc^{1/2}Re^{1/2}.$ (9)
At early times, the evaporative term in Eq. (5) is always negligible compared to the capillary pressure drainage term ($\tau_{d}/\tau_{e}\approx 10^{-4}$, while the dimensionless thickness is of order one), allowing us to recover the classical drainage law for surface bubbles $\tilde{h}\propto\tilde{t}^{-2/3}$. Evaporation becomes the main draining mechanism when the third term becomes of the same order of magnitude as the second one i.e. $\tilde{h}^{5/2}\sim\tau_{d}/\tau_{e}Sh(1-\mathcal{R}_{h})$. Using the early time scaling of the thickness over time $\tilde{h}\propto\tilde{t}^{-2/3}$, this gives the typical dimensionless time $t_{c}/\tau_{d}$ when evaporation becomes the main draining mechanism:
$t_{c}/\tau_{d}=\left(\frac{\tau_{d}}{\tau_{e}}Sh(1-\mathcal{R}_{H})\right)^{-3/5}.$ (10)
Figure 5: Thickness as function of dimensionless time. Markers indicate relative humidity and color shows the dimensionless evaporative rate. Full lines are solutions of Eq. (5), with $\alpha_{d}=0.5$ and $\alpha_{e}=30$. The data shown was obtained with a solution of 200 µmol/L of SDS.
We validate this model only taking into account wind speed through evaporation using experimental data. We analyze the effect of humidity and wind on bubble drainage by measuring the thickness of the cap over time. To do so, we use a high speed camera to record the bursting event. This allows us to track the retraction of the rim as a hole opens in the cap. The thickness is then accessible using the classical Taylor-Culick retraction velocity: $V_{\textsc{TC}}=\sqrt{2\gamma/(\rho h)}$ (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018; Shaw and Deike,, 2024). We repeat this process for several humidity values and for about 20 bubbles each time. We estimated the thickness at generation by using an average over the earliest bursting events we recorded (with a lifetime shorter than 1 second), and found this initial thickness to be independent of wind speed and humidity and equal to $h_{0}=12$ µm.
The result is shown in Fig 5 using an SDS solution (200 µmol/L): the thickness is normalized by the initial thickness $h_{0}$ and the time is normalized using $\tau_{d}$. For $t/\tau_{d}\lesssim 100$, we find that all the data for different evaporative rates (humidity or wind speed) collapse on the same trend, close to $h\propto t^{-2/3}$. Finding that this classical result is unchanged shows that the strong mixing of the cap at high wind speed actually has little effect on the drainage of the bubbles. Further, the motion of the marginal regeneration patches (their rise velocity or the size at which they detach from the foot of the bubble) does not seem to have a noticeable effect on the drainage in this case.
After a time which is typically of order $t_{c}$ (or $t/\tau_{d}\approx 100$), we see the cases at different evaporative rates separate and the drainage becomes faster. The thickness now decreases linearly with time, with the cases with strong evaporation (yellow) decreasing faster and earlier than the cases with little to no evaporation (black). We can quantitatively predict this behavior by integration of Eq. (5) over time combined with the Sherwood number given by Eq. (9) by fitting the numerical prefactors $\alpha_{d}$ and $\alpha_{e}$ on the data. We see a good agreement with our measurements, further validating the fact that the main draining mechanisms are i) viscous limited flow through the foot of the bubble and ii) direct evaporation from the cap. The numerical value of the prefactors we obtained are $\alpha_{d}=0.5$ and $\alpha_{e}=30$ for the whole dataset. The numerical integration slightly overshoots the cases at high humidity and low wind, possibly because of the hypothesis made in estimating the natural convection Sherwood number.
4.3 Lifetime of solutions of pure surfactants
Figure 6: (a) Mean lifetime as a function of the inverse of the magnitude of the evaporative rate. (b) Dimensionless lifetime as a function of the time $t_{c}$ (defined in Eq. (10)) when evaporation becomes the main draining mechanism. Markers show relative humidity and colors show SDS concentration. (orange line) Numerical integration of Eq. (5), with $\alpha_{d}=0.5$ and $\alpha_{e}=12.5$. (black dashed line) Scaling law $t_{f}=0.4t_{c}$.
In this section, we analyze the lifetime and drainage of surface bubbles in a solution only containing SDS (no salts). Plotting our data for several wind speeds, relative humidities and SDS concentrations (Fig. 6 (a)) reveals that most of the changes in mean lifetime are captured using a single parameter that contains the effects of wind and humidity on the evaporative rate: $\left(Sh\times(1-\mathcal{R}_{H})\right)^{-1}$. This result confirms that our modelling of the Sherwood number in the large range of conditions (from quiescent to fully turbulent) is correct. In addition, it shows that there is no visible effect of the wind on bubble lifetime beyond altering the evaporative rate. In this case with no salts in the solution, the effects of the surfactant concentration are very small and do not show a significant trend. Under the conditions explored in this study, the mean lifetime of bubbles in solutions containing only deionized water and surfactants is always larger than about 10 seconds (even though individual bubbles still sometimes burst earlier than 10s, see Fig 4 (b)).
We expect the lifetime of these bubbles to be of the same order of magnitude as the time $t_{c}$ defined in Eq. 10, when evaporation becomes the main draining mechanism. To test our scaling, we plot the same lifetime data as a function of $t_{c}$ (both times normalized by $\tau_{d}$) in Fig. 6 (b). This rescaling indeed captures well the evolution of the lifetime as $t_{f}/\tau_{d}$ grows linearly with $t_{c}/\tau_{d}$ with a slope of 0.4 (dashed line). Most of the spread in the data at high wind speed (lower left corner) that was present in Fig. 6 (a) is also captured by using $t_{c}/\tau_{d}$ showing that it was most likely due to small variations in surface tension or bubble size.
The same result can also be obtained numerically: we set a fixed thickness bursting criterion at a very small thickness ($h_{c}=10^{-3}h_{0}$), integrate Eq. (5) until the thickness reaches $h_{c}$ and record the associated time. This allows us to obtain the orange line, that also follows $t_{f}\sim t_{c}$ (shown analytically if $h_{c}=0$ in (Poulain and Bourouiba,, 2018)) and predicts the data very well if the prefactor $\alpha_{e}$ is slightly adjusted compared to its value obtained in the drainage experiments (here $\alpha_{d}=0.5$ and $\alpha_{e}=12.5$). This change is most likely due to small experimental differences between the drainage and lifetime setups: the tank in which the drainage experiment is performed is smaller such that the bubble always remains in view of the high speed camera. The evaporative prefactor $\alpha_{e}$ being larger in that case can therefore be interpreted as the influence of the proximity of the bath on the evaporative rate being less important.
Poulain and Bourouiba, (2018) observed for bubbles in the large surfactant concentration limit, that the longest time a surface bubble can live is given by $t_{c}$, when evaporation becomes the dominant drainage mechanism. With our setup in which we took great care in removing impurities (see Sec 3), we find that the average lifetime is of order $t_{c}$ (instead of just the maximum). This result is valid for very large changes of SDS concentration, from values greater than the CMC down to just a few µmol/L (or $10^{-3}$ times the CMC). All of our lifetime data for SDS is well collapsed when plotted as a function of $t_{c}$. This result is particularly robust at high wind speeds (or low $t_{c}$), which is coherent with our previous observations that mixing the air around the bath and the bubble improves the repeatability of lifetime experiment by preventing local fluctuations of the humidity field. At high $t_{c}$, we observe lifetimes that can be larger than the trend $t_{f}/\tau_{d}=0.4\>t_{c}/\tau_{d}$, and in general more important fluctuations. These may be due to the difficulty in modeling the Sherwood number associated with natural convection.
4.4 Lifetime of solutions containing salts
The lifetime of bubbles in solutions of salt and surfactants is drastically different from that of bubbles in pure water and surfactants. The lifetime data is summed up in Fig. 7 for sea salt (a mixture of salts and inorganic compounds in their oceanic proportions) and NaCl only (35 g/L). The top row shows the dimensional lifetime as a function of the inverse of the magnitude of the evaporative rate $\left(Sh\times(1-\mathcal{R}_{H})\right)^{-1}$, illustrating that in this case too, this parameter is adequate to combine the effects of changing wind speeds and relative humidities. The bottom row shows the same data, made dimensionless in the same way as in Fig. 6, using the drainage timescale $\tau_{d}$. In contrast to the previous case where concentration did not have a large effect and all bubble burst at a time given by $t_{c}$, here we see in both cases a transition between a regime where bubble lifetime is much shorter than $t_{c}$ and independent of relative humidity or wind speed, and a regime again dominated by evaporation, with the lifetime given by Eq. (10), with the same prefactor as in Fig. 6 for pure water and SDS.
At low concentrations ($c_{\textsc{SDS}}\lesssim 50$ µmol/L), the lifetime is essentially independent of wind speed or relative humidity. In this regime, the lifetime of solutions with sea salt is about 5 s and 10 s with NaCl. It is interesting to note that in both cases in this limit the lifetime is again independent on the surfactant concentration within the range tested. Even though the concentrations in SDS are fairly small (less than 1% of the CMC), the combined presence of salts and surfactants results in a large variation of the surface tension (see Appendix A).
Figure 7: (a) and (b) Lifetime as function of inverse of the magnitude of the evaporative rate for the sea salt and NaCl cases, respectively. (c) and (d) Dimensionless lifetime as a function of the dimensionless time when evaporation becomes dominant $t_{c}/\tau_{d}$. The left column shows the dataset with sea salt and the right one with NaCl. The dashed lines are identical to Fig. 6 (c): linear trends with a slope of 0.4. The orange lines are obtained by integrating of Eq. (5) until the thickness reaches $h_{c}$ (with the same prefactors as in Fig. 6). Markers show relative humidity and color shows SDS concentration.
Increasing the SDS concentration from this limit, we see the lifetime in both cases start to increase and to show a positive trend with the timescale $t_{c}$. In intermediate cases (100 and 200 µmol/L) the increase of lifetime with $t_{c}$ is slower than predicted by Eq. (10) and the lifetime always remains inferior to the pure water and SDS case ($t_{f}/\tau_{d}=0.4~{}t_{c}/\tau_{d}$, dashed line). For the highest concentrations, the lifetime follows this trend again, showing that the drainage has reached an evaporation dominated regime. The longest lifetimes obtained are similar for the cases with pure water or either salts and equal to about one minute. The case with NaCl and the largest amount of surfactant is noticeably above the rest of the data points, following the same linear trend but with a prefactor close to 0.6. It should be noted that in both cases this regime corresponds to an almost saturated isotherm (with a very low surface tension that does not vary with surface compression), similar to what would be obtained for a solution above the CMC. Finally, in the case of sea salt, the lifetime data departs from the trend at higher $t_{c}$ (high relative humidity or low wind speed) even at the highest SDS concentration, and increases slowly with $t_{c}$.
The behavior observed for surface bubbles with salts can qualitatively be replicated numerically by integrating Eq. (5) until the thickness reaches a finite thickness $h_{c}$. Results of this integration for various $h_{c}$ are shown in Fig. 7 in shades of orange, with the same numerical prefactors as in the case of SDS only ($\alpha_{d}=0.5$ and $\alpha_{e}=12.5$). With decreasing $h_{c}$ we see a transition from a regime where $t_{f}<t_{c}$ and $t_{f}$ is independent of relative humidity or wind speed to the same regime described in the case of pure SDS where $t_{f}\sim t_{c}$. This behavior qualitatively matches the one observed for sea salt or NaCl, showing that taking into account the effect of wind and relative humidity through the evaporative rate, and the effect of the solution through an average bursting thickness is enough to describe most of the trends seen in this study. The model predicts that the average bursting thickness in the wind independent regime is $h_{c}/h_{0}\approx 0.05$ or $h_{c}\approx 0.6$ µm for sea salt and $h_{c}/h_{0}\approx 0.02$ or $h_{c}\approx 0.2$ µm for NaCl. These values are coherent with what would be expected from bubbles with a radius of 1.6 mm with sea salt, extrapolating the data from (Shaw and Deike,, 2024) to our smaller size. Both cases become evaporation dominated when $h_{c}$ is less than 100 nm. Note that we have made here the hypothesis that the drainage curves are not altered by the presence of salts: Shaw and Deike, (2024) have recently measured the drainage of surface bubbles in a large range of conditions and have shown in particular that the effect of sea salt on drainage is much smaller than other effects such as solution temperature for instance and that the evolution of the thickness is still well captured by $h\sim t^{-2/3}$ at early times.
5 Discussion
The data presented in this article illustrates two very different regimes of surface bubble lifetime. For pure water and SDS only, the lifetime is essentially independent of the SDS concentration (from about one thousandth of the CMC to above the CMC) and closely follows Eq. (10). When either NaCl or sea salt is added, the lifetime is greatly reduced for a given SDS concentration and shows a transition from being independent of wind speed at low SDS concentrations to being dominated by evaporation at high concentrations as in the case of pure water and SDS.
We believe the difference between these two regimes lies in the presence of impurities in the solution, which are required to trigger a bursting event at a large thickness. A large thickness means here bubble caps thicker than 1 µm, in opposition to the very small thicknesses ($h\lesssim 0.1$ µm) attainable for long-lasting bubbles. A bubble of pure water and surfactant does not burst before evaporation becomes the dominant drainage mechanism and rapidly (linearly in time) removes material from the cap. The regime where evaporation is dominant has previously been identified as the upper limit on bubble lifetime (Poulain and Bourouiba,, 2018), the difference with our study is the use of the overflow to remove most co-surfactants from the interface and ensure a precise control of surface properties. Overflow seems an efficient way of also removing the small amount of impurities (other surface active compounds than the main surfactant) that may be present when using only deionized water and surfactants in addition to ensuring that to physicochemical properties of solution remain constant (see Sec. 3).
Authors have previously hypothesized that surface tension inhomogeneities (for instance across the marginal regeneration patches) may be enough to trigger bursting of bubble cap (Lhuissier and Villermaux,, 2012). From our data it seems that with SDS only, these inhomogeneities are not enough to make the bubble burst. This fact is supported by that wind, which drastically alters the marginal regeneration patches and presumably any Marangoni flow that might be present within the bubble cap, does not seem to have an effect on lifetime or drainage beyond a change of the evaporative rate.
Introducing salts changes the surface tension but also adds impurities to the solution. These impurities may then trigger large thickness bursting as theorized in Poulain et al., (2018) (Sec 6). The sea salt that we used contains many different salts but also other inorganic compounds and therefore more impurities than NaCl (which is 99% pure). This fact may help explain why for a given SDS concentration, the sea salt case always has a shorter lifetime than the NaCl case (5 versus 10 s in the lower SDS concentration limit). Finally, another clue towards the need for impurities to trigger bursting at a large thickness is a few experiments we performed with tap water instead of deionized water (therefore also containing impurities). The resulting mean lifetime was close to 5 s with a small concentration of SDS (10 µmol/L) in contrast to about 15 s with deionized water. We also performed experiments with Triton X-100, a non-ionic, slow adsorbing surfactant, which confirms the trends observed with SDS: for a given small concentration, the mean lifetime follows $t_{f}\sim t_{c}$ with surfactant alone, but it is a constant at around 5s when sea salt is added. These two test cases point towards impurities being the reason for the difference between the cases with and without salts, instead of an interaction specific to SDS and ions in solutions.
Knowledge of the Gibbs isotherm of the solution considered is not enough to predict the expected lifetime of a surface bubble. Indeed, the amount of impurities cannot be directly inferred from the isotherm and therefore neither can the expected bursting thickness. A method of measuring the nature and concentration of impurities in a solution may therefore help in predicting the exact lifetime of a surface bubble.
Acknowledgement This work was supported by NSF grant 2242512, NSF CAREER 1844932 and the Cooperative Institute for Modeling the Earth’s System at Princeton University to L.D.
Declaration of Interests. The authors report no conflict of interest.
Appendix A Langmuir trough measurements
Every solution used in this study was tested in a Langmuir trough (KSV NIMA, model KN 1003) using the Wilhelmy plate method to record its surface tension isotherm. We record the properties of the surface by setting aside a small amount of solution and placing it in the trough, and the surface tension is measured with a platinum Wilhelmy plate. After that, two Teflon barriers compress the surface at a rate of 270 mm/min allowing us to obtain the surface tension $\gamma$ as a function of the through area. We performed the measurement as quickly as possible after pouring the solution into the trough to avoid the adsorption of co-surfactants to the interface. The resulting data is plotted in Fig. 8 as a function of the area, normalized by the initial area of the trough.
The range of concentrations tested with SDS only corresponds to the full range of interface properties from very clean (flat isotherm at 72 mN/m) to an almost fully surfactant packed interface (flat isotherm at 30 mN/m). The isotherms with salts also show a broad range of interface properties, but with a much narrower concentration range. Because of other surfactants present in the salts used and of the interaction between SDS and ions in the solution the maximal surface tension obtained in those cases is 63 mN/m with NaCl and 54 mN/m with sea salt. Both of these cases reach a fully packed case at 400 µmol/L of SDS with a flat isotherm at about 30 mN/m.
Figure 8: Gibbs isotherms of the various solutions used in this study. Numbers indicate the SDS concentration used in µmol/L.